OpenStudy (abhisar):

An electron is moving at \(\sf 10^6\) m/s in a direction parallel to a current of 5A flowing through an infinite long straight wire separated by a perpendicular distance of 0.1 m in air. Magnetic force experienced by the electron will be?

OpenStudy (abhisar):

This is how I am doing.... The field due to the infinite current wire at any point at a perpendicular distance of 0.1m from it can be calculated by ampere's law as, \(\sf B=\Large \frac{5 \times 2\mu_0}{4\pi\times 0.1}=10^2\times 10^{-7}=10^{-5}\) Now, magnetic force can be calculated as, \(\sf \vec{F_m}=q(\vec{V}\times \vec{B})\\ \Rightarrow 1.6\times 10^{-19}(10^6\times 10^{-5}\times sin(90^0))=1.6 \times 10^{-18}\)

OpenStudy (rushwr):

Ur methods are correct I suppose ! Still having a slight doubt on the 1st part . Applying of F=BqV is alright. What's the answer given ?

OpenStudy (abhisar):

Answer is different doe, the key says \(\sf 4 \times 10^{-11}\) N :(

OpenStudy (abhisar):

Whoops... Wait !! The answer is \(\sf 1.6 \times 10^{-18}\)N. All this time I was looking answer for different question :D.

OpenStudy (rushwr):

lol

OpenStudy (rushwr):

ok then

OpenStudy (abhisar):

Thnx doe c:

OpenStudy (rushwr):

lol no problem